OpenStudy (anonymous):

A rectangle has an area of 12 square centimeters and a perimeter of 16 centimeters. Which of the following could be its dimensions

OpenStudy (anonymous):

xy=12 2(x+y)=16 x+y=8 x=8-y (8-y)y=12 y^2-8y+12=0 y^2-6y-2y+12=0 (y-6)(y-2)=0 y= 6 y=2 x=2 x=6